OpenStudy (anonymous):

can anyone help me with this and show how you got the answer e^(.05t)=1600

OpenStudy (anonymous):

if you want to get rid of e you need to apply "ln" to both sides of your problem. this will leave you with (0.05t)=ln(1600). From here it is a one step division problem.

jimthompson5910 (jim_thompson5910):

\[\large e^{0.05t}=1600\] \[\large \ln(e^{0.05t})=\ln(1600)\] \[\large 0.05t\ln(e)=\ln(1600)\] \[\large 0.05t(1)=\ln(1600)\] \[\large 0.05t=\ln(1600)\] \[\large t=\frac{\ln(1600)}{0.05}\]

OpenStudy (anonymous):

ln(e)?

OpenStudy (anonymous):

\[e^{0.05t}=1600\iff 0.05t=\ln(1600)\] equivalent logarithmic form

OpenStudy (anonymous):

just thought i would mention it

jimthompson5910 (jim_thompson5910):

I pulled down the exponent and evaluated the natural log of e to get 1 (ie ln(e) = 1)

OpenStudy (anonymous):

so is the answer going to be t=20ln(1600)

OpenStudy (anonymous):

yeah i see and of course it is right, but \[e^{x}=y\] and \[x=\ln(y)\] are the same statement therefore \[e^{0.05t}=1600\iff 0.05t=\ln(1600)\]

jimthompson5910 (jim_thompson5910):

yes, the answer is \[\large t=20\ln(1600)\]

OpenStudy (anonymous):

Ok thanks last assignment for my math before I am done for good with it .